The Law School Admission Test (LSAT) is required for admission to many law schools. The scores on the LSAT are approximately normal with mean of 150 and standard deviation of 9. Queenís School of Business in Kingston, Ontario requires a minimum LSAT score of 157 for admission.

1. a)What is the z-score corresponding to an LSAT score of 157? Round your answer to three decimal places.
z = .778

b)What is the probability that a randomly selected person who took the LSAT would score above 157? Round your answer to three decimal places.
Answer= .218



Bellow are the questions i don't understand!!!

c)What is the probability that the mean score for 5 randomly selected people who took the LSAT would be above 157? Round your answer to three decimal places.

d)What is the probability that the mean score for 10 randomly selected people who took the LSAT would be above 157? Round your answer to three decimal places.

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

For parts c) and d), you would use the formula SE = s/sqrt(n) where


SE = standard error (this is the standard deviation of the xbar distribution)
s = standard deviation (in this case, s = 9)
n = sample size (for part c), n = 5 ... for part d), n = 10)

-------------------------------------------------------

Once you've computed the standard error (SE), you can now recompute the z-score.

You will use the formula

z = (x-mu)/SE

and then you find the area to the right of that z-score using a calculator or table.

Answer by stanbon(75887)

The Law School Admission Test (LSAT) is required for admission to many law schools. The scores on the LSAT are approximately normal with mean of 150 and standard deviation of 9. Queenís School of Business in Kingston, Ontario requires a minimum LSAT score of 157 for admission.
---------------------------------------------------
1. a)What is the z-score corresponding to an LSAT score of 157? Round your answer to three decimal places.
z = .778
---------------------------------------------------
b)What is the probability that a randomly selected person who took the LSAT would score above 157? Round your answer to three decimal places.
Answer= .218
---------------------------------------------------
Bellow are the questions i don't understand!!!
c)What is the probability that the mean score for 5 randomly selected people who took the LSAT would be above 157? Round your answer to three decimal places.
The standard deviation is 9/sqrt(5)
z(157) = (157-150)/(9/sqrt(5)) = (7/9)*sqrt(5) = 0.778*sqrt(5) = 1.739
==================================
d)What is the probability that the mean score for 10 randomly selected people who took the LSAT would be above 157? Round your answer to three decimal places.
Ans: (0.778)*sqrt(10) = 2.46
=====================================
Note:: Check the Central Limit Theorem to under stand why
the standard deviation changes when you have a sample of data.
-----------------------------------------
Cheers,
Stan H.
----------------------